Solve the following system of equations. Please show your work to receive full credit.
x - y = 10
2x + y = 2
Explain which method you chose to solve the system and why you felt it was the best choice.

Answer by mananth(16946)

x - y = 10 .............1
Total value
2 x + y = 2 .............2
Since the co-efficients of y are equal and have opposite signs, we choose to eliminate y
Eliminate y
1 x -1 y = 10
2 x + 1 y = 2
Add the two equations
3 x = 12
/ 3
x = 4
plug value of x in (1)
1 x -1 y = 10
4 -1 y = 10
-1 y = 10 -4
-1 y = 6
y = -6